OpenStudy (anonymous):

Probability: 15 new students are to be distributed evenly among three classes. There are 3 girls among the new students. a) What is the probability all 3 girls are assigned to the same class? b) What is the probability that each class gets one girl? (Work to follow)

OpenStudy (anonymous):

For a, multinomial: Number of ways students can be assigned without restriction: \[\left[\begin{matrix} & 15 & \\ 3 & 3 & 3\end{matrix}\right]\] = 756,756

OpenStudy (anonymous):

There are 3 classes to assign, so \[3*\left(\begin{matrix}12 \\ 2\end{matrix}\right) \left(\begin{matrix}10 \\ 5\end{matrix}\right)\left(\begin{matrix}5 \\ 5\end{matrix}\right)\] = 49,896 ways to assign the three girls to one class room, therefore P(all girls assigned to same class) = 49,896/156,156 = .0659

OpenStudy (anonymous):

For b) there are \[\left(\begin{matrix}3 \\ 1\end{matrix}\right)\left(\begin{matrix}12 \\ 4\end{matrix}\right)\left(\begin{matrix}2 \\ 1\end{matrix}\right)\left(\begin{matrix}8 \\ 4\end{matrix}\right)\left(\begin{matrix}1 \\ 1\end{matrix}\right)\left(\begin{matrix}4 \\ 4 \end{matrix}\right)= 3*495*2*70*1*1= 207,900\] ways to distribute 1 girl to each class. Thus P(one girl to each new class) = 207900/756,756 = .2747 Did I miss anything?

OpenStudy (amistre64):

I wish I could tell if you did or not :) i take it this is past elementary statistics

OpenStudy (anonymous):

It's the first week of probability. ;) Thanks anyway.

OpenStudy (anonymous):

ok i got something different for first one but i didn't use multinomial. i did this

OpenStudy (anonymous):

lets assume all girls get assigned to room 1 probability is \[\frac{3}{15}\times \frac{2}{14}\times\frac{1}{13}\]

OpenStudy (anonymous):

so multiply this by 3 because the events "all girls get assigned to room one" and "all girls get assigned to room 2" and "all girls assigned to room 3" are disjoint

OpenStudy (anonymous):

so i get \[\frac{18}{2730}=.00659\] rounded

OpenStudy (anonymous):

which looks suspiciously like what you got but i think my decimal is correct and you are off by one place

OpenStudy (anonymous):

Hmm. That is interesting....I will play around with it more. Did you agree with part b?

OpenStudy (anonymous):

second one i think is right but i would do it differently. i think you just need to use your calculator again for first one. i bet we got same answer

OpenStudy (anonymous):

I did...I realize I typed the denominator incorrectly (should be 756,756) but I still get .0659.

OpenStudy (anonymous):

first of all \[\dbinom{1}{1}=\dbinom{4}{4}=1\] so forget the last terms

OpenStudy (anonymous):

sell maybe your denominator is off.

OpenStudy (anonymous):

Where did you get 18 in part a?

OpenStudy (anonymous):

lets compute the probability that 1 girl gets assigned to room 1. it is \[\frac{3}{15}\times \frac{12}{14}\times \frac{11}{12}\] and then the probability that one girl is assigned to room 2 given that one girl was assigned to room one is \[\frac{2}{11}\times \frac{10}{10}\times \frac{9}{9}=\frac{2}{11}\]

OpenStudy (anonymous):

oh i got 18 because i computed the probability that 3 girls got in room 1. that probability is \[\frac{6}{2730}\] but there are three rooms so i multiplied by 3 to get \[\frac{18}{2730}\]

OpenStudy (anonymous):

i get to multiply by three because as i said the events are disjoint

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok and finally if there is one girl in room 1 and one girl in room 2 there will be one girl in room 3 with probability 1. so my answer is \[\frac{3}{15}\times \frac{12}{14}\times\frac{11}{13}\times \frac{2}{11}\]

OpenStudy (anonymous):

let's see if it is the same as your answer. the problem with doing it with multinomial is that a huge compound fraction is forced on you so calculation is a bear. but if you do it without multiplying out you will see a plethora of cancellation and probably get exactly what i get

OpenStudy (anonymous):

This is a frustrating thing in Probability (to me), there are many different ways to get the same answer. Arg! I don't think we agree but I'm not sure why. Thanks for your help, feel free to move on if you'd like. ;)

OpenStudy (anonymous):

i get \[\frac{12}{455}=.02637\]\]

OpenStudy (anonymous):

oh crapola!

OpenStudy (anonymous):

i am so sorry

OpenStudy (anonymous):

i was putting 3 people in a class instead of 5!

OpenStudy (anonymous):

lets knock the first one out again quickly.

OpenStudy (anonymous):

Hahahaha awesome. Maybe that will help me out some cause I could not figure out what you were doing. Hehe.

OpenStudy (anonymous):

ok 3 girls in same class. class 1 that is. you get \[\frac{3}{15}\times \frac{2}{14}\times \frac{1}{13}\times 10\] damn you are right and i am wrong for sure sorry to waste your time, except i still think this is easier. then ten because there are ten ways to put 3 girls in 5 slots

OpenStudy (anonymous):

mutiply this by 3 and get exactly your answer because my decimal was off not yours

OpenStudy (anonymous):

it is not \[\frac{18}{2730}\] but rather \[\frac{180}{2730}=\frac{18}{273}=.0659\] what you said! so sorry

OpenStudy (anonymous):

That is helpful. So you are getting your fractions by choosing girls instead of classrooms, basically, right?

OpenStudy (anonymous):

yes. i put them all in one classroom and then multiply by 3

OpenStudy (anonymous):

but like an idiot i forgot there were 5 people in the class!

OpenStudy (anonymous):

so i had to multiply by ten, the number of ways to choose 3 from 5

OpenStudy (anonymous):

Awesome. definitely quicker/easier than multinomial. I like having a different method to check. Right... 5C3. Makes sense. Thank you for sticking with it. These can take so damn long.

OpenStudy (anonymous):

i am so sorry i forgot there were 5 people in a class and not 3. must have screwed you all up

OpenStudy (anonymous):

ok now lets do last one quickly. we put 1 girl in class 1

OpenStudy (anonymous):

Well, I wasn't convinced I was wrong since I couldn't talk myself into agreeing with your numbers so it wasn't too bad! ;)

OpenStudy (anonymous):

\[\frac{3}{15}\times \frac{12}{14}\times \frac{11}{13}\times \frac{10}{12}\times \frac{9}{11}\times 5\] and now we put 1 girl in class 2 given there is a girl in class 1 \[\frac{2}{10}\times \frac{8}{9}\times \frac{7}{8}\times \frac{6}{7}\times \frac{5}{6}\times 5\]

OpenStudy (anonymous):

we do a ton of cancellation and then multiply these together.

OpenStudy (anonymous):

aaaaand we agree, @ .2747. Sweet! Thanks bunches.

OpenStudy (anonymous):

after cancelling everything in sight i get \[\frac{5\times 5}{7\times 13}=\frac{25}{91}=.2747\]

OpenStudy (anonymous):

sorry i didn't just look at what you wrote, pretended i got it and said "yes you are right"

OpenStudy (anonymous):

ha! Thanks again for your help, I really appreciate it.

OpenStudy (anonymous):

i still like this way though. hate dealing with huge numbers in numerator and denominator. you are welcome

OpenStudy (anonymous):

Yes, it definitely simplifies some things. I always like having a new way to do something (especially when it saves time).